OpenStudy (anonymous):

out of 24 hours, how many hours is 20%. and how many is 31%? medal. thanks :)

OpenStudy (johnweldon1993):

so you would set up your equation like \[\frac{ x }{ 24 } = \frac{ 20 }{ 100 }\] then cross multiply....and solve for 'x' the same for 31% \[\frac{ x }{ 24 } = \frac{ 31 }{ 100 }\] hope that helps :)

OpenStudy (anonymous):

ohh, pellet.. thanks :) i remeber there was an easy way to do it, i just forgot. lol, sorry. thaniks soo much :) @johnweldon1993

OpenStudy (johnweldon1993):

haha you're welcome! :)